# The role of equation of state on the spin-up of millisecond pulsars

### v2

@@ 4 Comparison with observations
The recycling of pulsars is usual studied by comparison with observations in $P\dot{P}$ diagram. As we mentioned above, we keep $B$ as a parameter instead of replacing it by combination of $P$ and $\dot{P}$. Therefore, as shown in equations (1) and (2), in this paper the quantities to be compared with observations are $M$ and $P$.
If the values of both $B$ and $\Delta M$ of a recycled pulsar is known, we can plot the allowed region in $MP$ diagram which satisfies both equations (1) and (2), i.e. by putting Fig.1 and Fig.2 together. We show the allowed regions of SSs (left) and NSs (right) for $\Delta M=0.05M_{\odot}$ and $B=5\times 10^{8}$ G in Fig.3, where dashed line denotes the general spin-up line (cf. Fig.1) and solid line shows the minimal spin-period by accretion (cf. Fig.2). From this example we can see that, the allowed region of SSs covers the lower left region in $MP$ diagram where NSs cannot reach, indicating that it could be easier for an SS to be fully recycled than that for an NS with the same $B$ when the mass is below about $1.5M_{\odot}$.
Refer to captionRefer to caption
Figure 3: The allowed regions of SSs (left) and NSs (right) in $MP$ diagram, where solid curves denote minimal-spin lines with $\Delta M=0.05M_{\odot}$ and $B=5\times 10^{8}$ G, and dashed curves denote spin-up lines with $B=5\times 10^{8}$ G.
However, in most cases we are not certain about both the value of $B$ and $\Delta M$ for recycled pulsars. If we estimate $B$ by considering that, the vacuum magnetic dipole model with inclination angle $\alpha$ gives $B_{\rm dipole}\simeq 1.6\times 10^{9}\ {\rm G}\ \sqrt{P\dot{P}}\ \sin\alpha^{-1}$, then by setting $\sin\alpha=1$ we can get $B>1.6\times 10^{9}\ {\rm G}\ \sqrt{P\dot{P}}$. It is interesting to noted that, each pulsar whose data of $M$ and $P$ is plotted in Fig.5 (see below) is above the spin-up line of $B=1.6\times 10^{9}\ {\rm G}\ \sqrt{P\dot{P}}$ calculated using the data of $P$ and $\dot{P}$ from the ATNF Pulsar Catalogue (https://www.atnf.csiro.au/research/pulsar/psrcat/(Manchester et al., 2005) under both SS and NS models. As mentioned before, in this paper we keep $B$ as a parameter, instead of replacing in some model-dependent ways. Then how to show the difference between SSs and NSs models by comparing with observations?
Combining equations (1) and (2) we can get the minimal-spin contour by accretion of $\Delta M$, which is defined as the contour of $\Delta M$ in $MP$ diagram, representing the minimal spin-period by accretion of $\Delta M$ for different values of $B$. For a data point in $MP$ diagram, there exists both a minimal-spin contour and a spin-up line intersecting it. The value of $\Delta M$ of the minimal-spin contour is the minimal accreted mass of the pulsar, while the value of $B$ of the spin-up line is the upper limit on its magnetic field strength. If we can use some other methods to constrain $\Delta M$ (e.g. from the binary evolution) and $B$ (e.g. theoretically from the magnetic field decay in accretion, or observationally from the cyclotron absorption lines), it would be hopeful to see that whether this pulsar is more likely to be an NS or an SS.
From Fig.3 we can infer that low mass and fast spinning pulsars would present significant different parameters between SS and NS models, so we use a hypothetical data point in $MP$ diagram as an example in Fig.4. If this pulsar is an SS, a minimal-spin contour of $\Delta M=0.059M_{\odot}$ (red line) and a spin-up line of $B=6.6\times 10^{8}$ G (red dashed line) intersect it in $MP$ diagram, indicating that the accreted mass $\Delta M>0.059M_{\odot}$ and its magnetic field strength $B<6.6\times 10^{8}$ G. One the other hand, if this pulsar is an NS, a minimal-spin contour of $\Delta M=0.060M_{\odot}$ (blue line) and a spin-up line of $B=4.1\times 10^{8}$ G (blue dashed line) intersect it in $MP$ diagram, indicating that the accreted mass $\Delta M>0.060M_{\odot}$ and its magnetic field strength $B<4.1\times 10^{8}$ G. This example shows that the difference in the the upper limit on $B$ between SS and NS models is significant for low mass and fast spinning pulsars.
Figure 4: We use a hypothetical data point in $MP$ diagram as an example. If this pulsar is an SS, a minimal-spin contour of $\Delta M=0.059M_{\odot}$ (red line) and a spin-up line of $B=6.6\times 10^{8}$ G (red dashed line) intersect it, indicating that the accreted mass $\Delta M>0.059M_{\odot}$ and its magnetic field strength $B<6.6\times 10^{8}$ G. If this pulsar is an NS, a minimal-spin contour of $\Delta M=0.060M_{\odot}$ (blue line) and a spin-up line of $B=4.1\times 10^{8}$ G (blue dashed line) intersect it, indicating that the accreted mass $\Delta M>0.060M_{\odot}$ and its magnetic field strength $B<4.1\times 10^{8}$ G.
@@
To compare with actual data, we extend the curves in Fig.4 to different values of $B$ and $\Delta M$, and we show in Fig.5 the results for SSs (left) and NSs (right). Solid curves correspond to minimal-spin contours with $\Delta M/M_{\odot}=0.02,0.04,0.06,0.1,0.15,0.3$, increasing downwards, and dashed curves correspond to spin-up lines with $B_{\rm max}/10^{8}{\rm G}=2,4,8,12$, increasing upwards. To ensure that the data are of fully recycled pulsars, we choose the masses of pulsars whose $P<10$ ms from Özel & Freire (2016) and plot their $M$ (with 1-$\sigma$ uncertainties) and $P$ in this $MP$ diagram.
Refer to captionRefer to caption
@@
Figure 5: Minimal-spin contours of $\Delta M$ and spin-up lines of $B$ for SSs (left) and NSs (right). The horizontal ordinate denotes the final mass after accretion. Solid curves correspond to $\Delta M/M_{\odot}=0.02,0.04,0.06,0.1,0.15,0.3$, increasing downwards, and dashed curves correspond to contours for $B/10^{8}{\rm G}=2,4,8,12$, increasing upwards. Data points with 1-$\sigma$ uncertainties are from Özel & Freire (2016).
The different effects on parameters of recycled MSPs between SS and NS models can be seen from Fig.5. Although the minimal-spin contours of $\Delta M$ for SSs are not significantly different from that for NSs, the differences between the contours of $B$, i.e. the spin-up lines, for SSs and that for NSs are visible. The dashed curves in Fig. 5 show that the magnetic field strength will play an important role in impeding the spin-up of NSs, especially for the ones with low masses, but the impeding effect would not be important to SSs. The constraints on EoS models are more significant for masses below about $1.5M_{\odot}$ than that for higher masses.
@@
If we can find some fast spinning pulsars with low masses, and if we can additionally know the lower limit of $B$, i.e. $B>B_{0}$, it would be hopeful to see that whether these pulsars is more likely to be an NS or an SS. For example, if the magnetic field is higher than $4\times 10^{8}$ G ($2\times 10^{8}$ G), then an NS with $M<1.6M_{\odot}$ ($M<1.2M_{\odot}$) could not be spun up to $P<3$ ms ($P<2$ ms), but an SS could. As we mentioned before, each pulsar plotted in Fig.5 is above its spin-up line of $B=1.6\times 10^{9}\ {\rm G}\ \sqrt{P\dot{P}}$ calculated using its data of $P$ and $\dot{P}$, so at this point if we take the lower limit magnetic field strength $B_{0}$ as $1.6\times 10^{9}\ {\rm G}\ \sqrt{P\dot{P}}$ we could not yet disfavor neither SS nor NS models. Certainly, finding more MSPs of $P<10$ ms with low mass and more accurate estimation of $B_{0}$ in the future could help to put more strict constraints on the EoS of pulsars.
@@ 5 Conclusions and discussions
The global properties of pulsars determined by the EoS, such as mass, radius and the moment of inertia, could play a role in the spin-up process of accreting pulsars in binaries. We investigate the spin-up of MSPs in AP4 EoS of NS model and LX3630 EoS of the SS model. Combining the spin-up condition and the transferred angular momentum, we can constrain the accreted mass $\Delta M$ and the magnetic field strength $B$ of recycled MSPs. Under the same magnetic field strength, an SS could spin faster than an NS in the recycling process, and NSs would spin faster than SSs with the same amount of accreted mass. The minimal-spin contours show that to get the minimal spin the amount of accreted mass needed by an SS and an NS would be almost the same, but the differences between the spin-up lines for SSs and that for NSs are obvious. The magnetic field strength will play an important role in impeding the spin-up of NSs, especially for the ones with low masses, but the impeding effect would not be important to SSs, so it could be easier for an SS to form a low-mass and fully recycled MSP than that for an NS.
The constraints on EoS models are more significant for masses below about $1.5M_{\odot}$, when the upper limit of the allowed magnetic field strength of an SS is higher than that of an NS. To give a more general and model-independent investigation, in this paper we keep the magnetic field strength $B$ as a parameter instead of replacing it in some model-dependent ways. It is worth noting that an accreting pulsar would have an lower limit of magnetic field strength (e.g. see Zhang & Kojima (2006)), so the constraints on NS model would be more strict than that on SS model. For the fast spinning pulsars with low masses, if we can additionally derive the lower limit of magnetic field strength theoretically from the magnetic field decay models, or if we can directly detect the magnetic field strength from the cyclotron absorption lines, it would be hopeful to see that whether these pulsars is more likely to be an NS or an SS.
@@
Although there are a lot of work about the recycling of MSPs which mainly focus on the effects of the interactions between pulsars and donor stars during binary evolution, the mass and radius are usually fixed to be typical values, such as $1.4M_{\odot}$ and 10 km. We show in this paper, however, that different global structures could bring observational differences, implying that the effects brought by EoS models should be taken into account in studying the recycling of MSPs. To focus on the EoS-dependent results, it is inevitable to ignore the dependence on other factors in this paper, such as the pattern of disk-fed accretion flow and the magnetic field configuration outside the accreting star, which lead some simplifications in equations (1) and (2). In addition we do not consider the nature of the donor star relating to the amount of accreted mass and evolution of the magnetic field due to accretion. Still, our results indicate that it is worth adding the role of EoS to other effects to study the recycling of MSPs more comprehensively.
@@
The AP4 EoS and the LX3630 EoS respectively represent EoS models of gravity-bound and self-bound, so we take them as examples to derive some universal conclusions. A gravity-bound pulsar, as shown by AP4 EoS, would have nearly unchanged radius when the mass increases from 1 to 2 $M_{\odot}$, whereas the mass of a self-bound pulsar, as shown by LX3630 EoS, is nearly proportional to the third power of radius when the mass is well below the maximum value. This difference leads to different evolutions of $r_{\rm A}$ during mass accretion, resulting in different allowed regions in $MP$ diagram of recycled pulsars. We could make a general conclusion that, it is easier for a self-bound star to form a low-mass and fully recycled MSP than that for a gravity bound star. Although the we could not yet rule out neither SS nor NS models by the present data, finding more samples of fully recycled low-mass MSPs with accurate mass-measurement and better constraints on magnetic field strength in the future, e.g. by SKA and FAST, could be helpful to put more strict constraints on the nature of pulsars.
This work is supported by the National SKA Program of China (No. 2020SKA0120300) and the Young Top-notch Talent Cultivation Program of Hubei Province.

### v3

@@ 4 Comparison with observations
The recycling of pulsars is usual studied by comparison with observations in $P\dot{P}$ diagram. As we mentioned above, we keep $B$ as a parameter instead of replacing it by combination of $P$ and $\dot{P}$ to derive the spin-up lines in $MP$ diagram, and take into account the evolution of $B$ in the spin-up process. Therefore, as shown in equations (1) and (2), in this paper the quantities to be compared with observations are $M$ and $P$.
The spin-up lines derived in Section 2 do not consider the supplied mass, and the spin-periods after accretion of $\Delta M$ derived in Section 3 do not consider the spin-up condition, so we can derive the spin-periods after accretion of $\Delta M$ by putting Fig.2 and Fig.3 together. In addition, because we consider the evolution of $B$ during accretion, the spin-up lines depending on the value of $B$ can be determined by the accreted mass.
In the upper panel of Fig.4, the orange solid line shows the spin-periods of SSs in LX3630 model with mass $M$ after accretion of $\Delta M=0.1M_{\odot}$, and the orange dashed line is the spin-up line of SSs in LX3630 whose magnetic field strength equals to the final $B$ after accretion of $\Delta M=0.1M_{\odot}$. Therefore, when $\Delta M=0.1M_{\odot}$, the spin-period of an SS in LX3630 lies in the curve formed by taking the higher parts of these two lines. The red solid and dashed lines are the corresponding results of SSs in Z-2023 model. From the observational point of view, it would be easier to constrain the upper limit of $\Delta M$, then we can get the allowed region which lies above both solid and dashed lines. We can see that the allowed region for SSs (covered by both orange and red patches) in LX3630 is bigger than and encloses that for Z-2023 (covered by red patch).
The lower panel of Fig.4 shows the same lines as in the upper one, for the results of NSs in AP4 model (blue solid and dashed lines) and X-2024 model (green solid and dashed lines). The allowed region for X-2024 model (covered by both green and blue patches) is bigger than and encloses that for AP4 model (covered by blue patch). Comparing the upper and lower panels of Fig.4 we can see that, the allowed region of SSs (in both LX3630 and Z-2023 models) covers the lower left region in $MP$ diagram where NSs (in both AP4 and X-2024 models) cannot reach, indicating that it could be easier for an SS to be spun-up to a period less than 3 ms than that for an NS, when the mass is below about $1.2M_{\odot}$.
Figure 4: The allowed regions of SSs (upper) and NSs (lower) in $MP$ diagram. The upper panel shows the spin-periods of SSs with mass $M$ after accretion of $\Delta M=0.1M_{\odot}$ in LX3630 (orange solid line) and Z-2023 (red solid line), and the spin-up line of SSs whose magnetic field strength equals to the final $B$ after accretion of $\Delta M=0.1M_{\odot}$ in LX3630 (orange dashed line) and Z-2023 (red dashed line). If $\Delta M<0.1M_{\odot}$, the allowed region for SSs in LX3630 is covered by both orange and red patches, and that for SSs in Z-2023 is covered by red patch. The lower panel of shows the same results as in the upper one, for NSs in AP4 model (blue solid line, blue dashed line, and blue patch) and in X-2024 model (green solid line, green dashed line, and patches in both blue and green).
As $\Delta M$ increases/decreases, all the lines for $P$ after accretion of $\Delta M$ (solid lines) and spin-up lines (dashed lines) in Fig.4 will shift downwards/upwards. The allowed region of each model will be above both the solid and dashed lines in the corresponding color. For some values of $\Delta M$, solid line will be totally above or below the dashed line. In all cases, for a certain value of $\Delta M$ the spin-period of an MSP lies in the curve formed by taking the higher parts of solid and dashed lines in the same color, and for an upper limit of $\Delta M$ the allowed region of an MSP lies above both solid and dashed lines in the same color.
The above example shows that, the role of EoS on the spin-up of MSPs could be manifested in low-mass (below about $1.2M_{\odot}$) case. As also shown in Fig.2 and Fig.3, the difference between SSs in LX3630 and NSs in AP4 is more pronounced than that between SSs in Z-2023 and NSs in X-2024 model.
@@
To compare with observational data, we extend the curves in Fig.4 to different values of $\Delta M$, and we show in Fig.5 the results for SSs (upper) and NSs (lower). Each line is derived by taking the higher parts of line of $P$ after accretion of $\Delta M$ and line of spin-up, for a certain value of $\Delta M$. In the upper panel, the results for SSs in SS model (LX3630) and hybrid SS model (Z-2023) are shown by red solid line and red dashed line, respectively. In the lower panel, the results for NSs in NS model (AP4) and hybrid NS model (X-2024) are shown by blue solid line and blue dashed line, respectively. In both panels, $\Delta M/M_{\odot}=0.04,0.06,0.1,0.2$, increasing downwards. We choose the masses of pulsars whose $P<15$ ms from Özel & Freire (2016) and plot their $M$ (with 1-$\sigma$ uncertainties) and $P$ in this $MP$ diagram.
@@
Figure 5: $M$-$P$ curves after accretion of $\Delta M$, for SSs in LX3630 (red solid line in upper panel), SSs in Z-2023 (red dashed line in the upper panel), NSs in AP4 (blue solid line in lower panel), and NSs in X-2024 (blue dashed line in lower panel). For each model, we show the results for $\Delta M/M_{\odot}=0.04,0.06,0.1,0.2$, increasing downwards. Data points with 1-$\sigma$ uncertainties are from Özel & Freire (2016).
The different effects on parameters of recycled MSPs between SSs and NSs can be seen from Fig.5. The discrepancy between SSs in LX3630 (red solid line) and NSs in AP4 (blue solid line) is bigger than that between SSs in Z-2023 (red dashed line) and NSs in X-2024 (blue dashed line). In the low-mass ($M$ below about $1.2M_{\odot}$) case, an SS can spin faster than an NS of the same mass by accreting the same amount of mass $\Delta M$, and the difference is more significant when $\Delta M<0.1M_{\odot}$.
@@
The constraints on EoS models could be better for masses below about $1.2M_{\odot}$ than that for higher masses. If we can find some fast spinning pulsars with low masses and additionally know the upper limit of $\Delta M$, it would be hopeful to see that whether these pulsars is more likely to be an NS or an SS. For example, if we can constrain that $\Delta M<0.04M_{\odot}$, then an NS with $M<1.2M_{\odot}$ could not be spun up to $P<11$ ms, but an SS could. Furthermore, we can see that if we only consider SSs in LX3630 model, the discrepancy between SSs and NSs would be more significant. Certainly, finding more MSPs of $P<15$ ms with low mass and more accurate estimation of $\Delta M$ in the future could help to put more strict constraints on the EoS of pulsars.
@@ 5 Conclusions and discussions
The global properties of pulsars determined by the EoS, such as mass, radius and the moment of inertia, could play a role in the spin-up process of accreting pulsars in binaries. We investigate the spin-up of MSPs of NSs (in AP4 and X-2024 models) and SSs (in LX3630 and Z-2023 models). Combining the spin-up condition and the transferred angular momentum, we can constrain the spin-period and mass of an MSP. When considering the accretion process, we take into account the evolution of magnetic field during accretion and the fastness-dependent torque. Our results show that, for the spin-up process of MSPs there could be observable differences between SSs and NSs.
The impeding effect of magnetic field on the spin-up of MSPs would be more significant for NSs than for SSs, especially for the ones with masses below about $1.4M_{\odot}$. By accreting the same amount of mass, an SS with mass below about $1.2M_{\odot}$ can spin faster than an NS of the same mass, implying that it could be easier for an SS to form a low-mass and fully recycled MSP than that for an NS. Therefore, our results show that the constraints on EoS models would be better in such low-mass case. For the fast spinning pulsars with low masses, if we can additionally constrain the amount of accreted mass $\Delta M$ (which could be inferred by the nature of donor stars), it would be hopeful to see that whether these pulsars is more likely to be an NS or an SS.
@@
Although there are a lot of work about the recycling of MSPs which mainly focus on the effects of the interactions between pulsars and donor stars during binary evolution, the mass and radius are usually fixed to be typical values, such as $1.4M_{\odot}$ and 10 km. We show in this paper, however, that different global structures could bring observational differences, implying that the effects brought by EoS models should be taken into account in studying the recycling of MSPs.
To focus on the EoS-dependent results, it is inevitable to ignore the dependence on other factors in this paper, such as $\phi$ determined by the pattern of disk-fed accretion flow and the critical fastness $\omega_{\rm c}$ determining the accretion torque. Different choices of $\phi$ and $\omega_{\rm c}$ will influence the quantitative results, shifting the curves in Figs.2, 3, 4, and 5 upwards or downwards, but the qualitative conclusion that an SS could be more easily to form a low-mass and fully recycled MSP than an NS will not change. In addition, the expressions of the fastness-dependent torque and the evolution of magnetic field during accretion are also model-dependent. Still, our results indicate that it is worth adding the role of EoS to other effects to study the recycling of MSPs more comprehensively.
The observable differences are manifested in the low-mass case, which can be inferred from the fact that the discrepancy of $M$-$R$ curves between the two SS models and the two NS models is more significant in such low-mass case. Following the same reasoning, the difference between SSs in LX3630 and NSs in AP4 is more pronounced than that between SSs in Z-2023 and NSs in X-2024. Although the differences between the two SS models and the two NS models are obvious enough, if we take LX3630 model and AP4 model to represent EoS models of gravity-bound and self-bound respectively, we could get more robust conclusions.
@@
A gravity-bound pulsar, e.g. shown by AP4 EoS, would have nearly unchanged radius when the mass increases from 1 to 2 $M_{\odot}$, whereas the mass of a self-bound pulsar, e.g. shown by LX3630 EoS, is nearly proportional to the third power of radius when the mass is well below the maximum value. This difference leads to different evolutions of $r_{\rm m}$ during mass accretion, resulting in different allowed regions in $MP$ diagram of recycled pulsars. We could make a general conclusion that, it is easier for a self-bound star to form a low-mass and fully recycled MSP than that for a gravity bound star. Although the we could not yet rule out neither SS nor NS models by the present data, finding more samples of fully recycled low-mass MSPs with accurate mass-measurement and better constraints on the amount of accreted mass in the future, e.g. by SKA and FAST, could be helpful to put more strict constraints on the nature of pulsars.
Acknowledgements
We would like to thank Prof. Renxin Xu in PKU, Prof. Wencong Chen in QUT, Prof. Chengjun Xia in YZU and Dr. Chen Zhang in HKUST for useful discussions, and to thank an anonymous referee for useful comments and suggestions. This work is supported by the National SKA Program of China (No. 2020SKA0120300).
Data Availability
The data underlying the work in this paper are available upon reasonable request.